# CC-GPX: Extracting High-Quality Annotated Geospatial Data from Common Crawl

### v1

@@ Abstract.
The Common Crawl (CC) corpus is the largest open web crawl dataset containing 9.5+ petabytes of data captured since 2008. The dataset is instrumental in training large language models, and as such it has been studied for (un)desirable content, and distilled for smaller, domain-specific datasets. However, to our knowledge, no research has been dedicated to using CC as a source of annotated geospatial data. In this paper, we introduce an efficient pipeline to extract annotated user-generated tracks from GPX files found in CC, and the resulting multimodal dataset with 1,416 pairings of human-written descriptions and MultiLineString vector data. The dataset can be used to study people’s outdoor activity patterns, the way people talk about their outdoor experiences, and for developing trajectory generation or track annotation models.
@@ 1. Introduction
In the past, researchers have studied the contents of Common Crawl for undesirable content (Luccioni and Viviano, 2021), extracted documents to create the largest Spanish-language crawling corpus (Gutiérrez-Fandiño et al., 2022), extracted parallel texts for machine translation (Smith et al., 2013), PDF documents (Turski et al., 2023) and Word documents with layout annotations (Weber et al., 2023), and built the popular C4 (the Colossal, Cleaned version of Common Crawl) dataset [^2] https://paperswithcode.com/dataset/c4 among many other things. However, to our knowledge, no research was dedicated to analysing the geospatial contents or extracting high-quality, annotated geospatial data from CC.
@@ 2. Data Collection and Processing
Hikers, cyclists, and in particular runners often track their activities using GPS-enabled mobile devices, such as Garmin smart watches (Karahanoğlu et al., 2021), and sometimes add detailed annotations describing their experience. Alternatively, people may plan their active journey beforehand using websites such as cycle.travel [^6] https://cycle.travel/, and add notes before or after the journey. In this paper, we decided to focus on the popular XML-based .gpx [^7] http://www.topografix.com/GPX/1/1/ (GPX Exchange Format) files that are most commonly used to share the activities, and ignore less popular formats such as .fit or .tcx which are very rare in CC.
@@ 2.5. Translating descriptions to English
The majority of the GPX files we identified are from European websites, and their descriptions are not in English. We use pycld2131313https://github.com/aboSamoor/pycld2 Python library to identify the original language. We remove tracks whose language is identified as Unknown as those often contain majority numbers or incoherent text, as well as all tracks whose descriptions are in languages that have 5 or fewer samples.
@@ 3. Dataset
The final dataset consists of 1,416 tracks (or around 1.5% of the initially deduplicated GPX files) and accompanying descriptions in both the original language and English. The examples of two tracks with descriptions are shown in Figures 1 and 2. The final dataset is a GeoPackage (.gpkg) file containing 17 properties described in Table 1. Code and final dataset are available on GitHub [^18] https://github.com/ilyankou/cc-gpx.
@@
The GPX files come from 135 unique domain names. The track descriptions come in 11 languages, with the majority being French (802), German (330), and English (110). The original descriptions are between 50 and 1999 characters long (for translated, between 34 and 1866), with a median of 303 characters (for translated, 285.5). The tracks originate in 25 countries, with most popular being France (710), Austria (164), Germany (133), Switzerland (116), and Italy (81). The shortest track is 607 m (1991 ft) while the longest is 99.3 km (61.7 mi). The average track length is 20.5 km (12.7 mi), while the median is much shorter at 12.1 km (7.5 mi). Histograms of route length (in km) and description length (in characters) are shown in Figure 3.
@@ 5. Conclusion
In this paper, we demonstrated that the CC corpus can be used as a novel source of annotated geospatial data. As a case study, we built an efficient pipeline to extract annotated outdoor activity tracks from GPX files that can be executed on a medium-powered laptop in a matter of hours. The resulting multimodal dataset contains textual descriptions of outdoor activities paired with MultiLineString vector geospatial data and can be used to study people’s outdoor activity habits, and the relationship between what people experience and what they say they experience. The dataset can be useful as a fine-tuning set for various GeoAI models, from trajectory to human-like description generators. We recognise that our case study captures just a fraction of all annotated geospatial data available in CC, and we intend to extend the scope in future work.

### v2

@@ Abstract.
The Common Crawl (CC) corpus is the largest open web crawl dataset containing 9.5+ petabytes of data captured since 2008. The dataset is instrumental in training large language models, and as such it has been studied for (un)desirable content, and distilled for smaller, domain-specific datasets. However, to our knowledge, no research has been dedicated to using CC as a source of annotated geospatial data. In this paper, we introduce an efficient pipeline to extract annotated user-generated tracks from GPX files found in CC, and the resulting multimodal dataset with 1,416 pairings of human-written descriptions and MultiLineString vector data from the 6 most recent CC releases. The dataset can be used to study people’s outdoor activity patterns, the way people talk about their outdoor experiences, and for developing trajectory generation or track annotation models. Our reproducible code is available on GitHub [^1] https://github.com/ilyankou/cc-gpx.
@@ 1. Introduction
In the past, researchers have studied the contents of Common Crawl for undesirable content (Luccioni and Viviano, 2021), extracted documents to create the largest Spanish-language crawling corpus (Gutiérrez-Fandiño et al., 2022), extracted parallel texts for machine translation (Smith et al., 2013), PDF documents (Turski et al., 2023) and Word documents with layout annotations (Weber et al., 2023), and built the ‘Colossal Clean Crawled Corpus’ (C4) dataset [^3] https://paperswithcode.com/dataset/c4 among other things. However, to our knowledge, no research was dedicated to analysing the geospatial contents or extracting annotated geospatial data from CC.
@@ 2. Data Collection and Processing
Hikers, cyclists, and in particular runners often track their activities using GPS-enabled mobile devices, such as Garmin smart watches (Karahanoğlu et al., 2021), and sometimes add detailed annotations describing their experience. Alternatively, people may plan their active journey beforehand using websites such as cycle.travel, and add notes before or after the journey. In this paper, we decided to focus on the popular XML-based .gpx [^7] http://www.topografix.com/GPX/1/1/ (GPX Exchange Format) files that are most commonly used to share the activities, and ignore less popular formats such as .fit or .tcx which are very rare in Common Crawl.
@@ 2.5. Translating descriptions to English
The majority of GPX files we identified are from European websites, and their descriptions are not in English. We use pycld2131313https://github.com/aboSamoor/pycld2 Python library to identify the original language. We remove tracks whose language is identified as Unknown as those often contain majority numbers or incoherent text, as well as all tracks whose descriptions are in languages that have 5 or fewer samples.
@@ 3. Dataset
The final dataset consists of 1,416 tracks (or around 1.5% of the initially deduplicated GPX files) and accompanying descriptions in both the original language and English, along with other properties described in Table 1. The examples of two tracks with descriptions are shown in Figures 1 and 2.
@@
The GPX files come from 135 unique domain names. Track descriptions come in 11 languages, with the majority being French (802), German (330), and English (110). The original descriptions are between 50 and 1999 characters long (for translated, between 34 and 1866), with a median of 303 characters (for translated, 285.5). The tracks originate in 25 countries, with most popular being France (710), Austria (164), Germany (133), Switzerland (116), and Italy (81). The shortest track is 607 m (1991 ft) while the longest is 99.3 km (61.7 mi). The average track length is 20.5 km (12.7 mi), while the median is much shorter at 12.1 km (7.5 mi). Histograms of route length (in km) and description length (in characters) are shown in Figure 3.
@@ 5. Conclusion
In this paper, we demonstrated that Common Crawl can be used as a novel source of annotated geospatial data. As a case study, we built an efficient pipeline to extract annotated outdoor activity tracks from GPX files that can be executed on a medium-powered laptop in a matter of hours. The resulting multimodal dataset contains textual descriptions of outdoor activities paired with MultiLineString vector geospatial data and can be used to study people’s outdoor activity habits, and the relationship between what people experience and what they say they experience. The dataset can be useful as a fine-tuning set for various GeoAI models, from trajectory to human-like description generators. We recognise that our case study captures just a fraction of all annotated geospatial data available in CC, and we intend to extend the scope in future work.